OpenStudy (anonymous):

The probability of drawing a blue marble from a bag is 2/5 . The probability of drawing a red marble from the bag is 1/5 . What is the probability of drawing a red marble or a blue marble from the bag? A.2/5 B.4/5 C.2/25 D.3/5

OpenStudy (anonymous):

Is it B?

OpenStudy (mathmate):

um...

OpenStudy (anonymous):

hmmm

OpenStudy (mathmate):

P(B)=2/5, P(R)=1/5 \(P(B\cap R)=P(B)+P(R)-P(B\cup R)\) But since red and blue are independent, \(P(B\cup R)=0\) so \(P(B\cap R)=P(B)+P(R)= ? \)

OpenStudy (mathmate):

Erratum: \(P(B\cup R)=P(B)+P(R)−P(B\cap R)\) But since red and blue are independent, \(P(B\cap R)=0\) so \(P(B\cup R)=P(B)+P(R)=?\)

OpenStudy (mathmate):

@VivaLaChloe1 Can you work out P(B)+P(R) and try again?

OpenStudy (anonymous):

So is it D? @satellite73

OpenStudy (anonymous):

@mathmate @tester97

OpenStudy (tester97):

I think its D :\

OpenStudy (anonymous):

Me too.

OpenStudy (mathmate):

Correct, it's D.

OpenStudy (anonymous):

add !

OpenStudy (anonymous):

@mathmate a slight typo when you wrote \[P(B\cup R)=0\]

OpenStudy (anonymous):

I AM AN IDIOT.I MEANT D NOT B

OpenStudy (anonymous):

Ugh i'm so blind

OpenStudy (mathmate):

@satellite73 Yes, it wasn't even a "slight" typo. lol Thank you for pointing it out. I got the cups and caps inverted. I put up an erratum below. Perhaps next time I just have to delete the mistaken one and start anew! :)